OpenStudy (anonymous):

what is the sine of -390 in fraction form and how would i find out how to do this on my TI-84 graphing calculator?

OpenStudy (anonymous):

Make sure the mode is set to degrees and not radians. Enter sin(-390 to find the decimal form. Then, press math and select >Frac. Press enter and you have your fraction

OpenStudy (anonymous):

Thank you so much! helped a lot!

OpenStudy (anonymous):

Tip: Don't depend on your calculator

OpenStudy (anonymous):

It /will/ screw you over.

OpenStudy (anonymous):

Yeah i have the basic points of the unit circle memorized but thats not one of them

OpenStudy (anonymous):

If you can, sketch theta. Remember that sin(k)=y/r, cos(k)=x/r, and tan(k)=y/x

OpenStudy (anonymous):

Oh okay, haven't tried that

OpenStudy (anonymous):

just to make sure I'm correct its (cos,sin) for how the ordered pairs on a nit circle are ordered right?